Consider 2 masses apart at rest in space far from any other object. As gravity pulls them together, then for each mass the


 a. speed increases and acceleration decreases.

 b. speed increases and acceleration is constant.

 c. speed increases and acceleration increases.

 d. speed decreases and acceleration is constant.
Answer #1

Option (c) is the correct answer.

Explanation -

Gravitational force od attraction between two objects of masses m1 and m2 is given as -

F = (G*m1*m2) / r^2

where. G = Universal gravitational constant.

r = distance between the two objects.

Now, as the objects will move closer, value of 'r' will decrease.

So, magnitude of F will increase.

acceleration, a = F/m will increase.

Also, with increase in time, t, magnitude of velocity will also increase. [v = u + a*t]

Therefore, option (c) is the correct answer.